OpenStudy (anonymous):

For the function f(x)=-2sin(4x-pi), find the amplitude, period, phase shift, right end point, x and y axis orientation, and vertical shift.

OpenStudy (anonymous):

This is what I have for: Amplitude=-2 Period=pi/2 Phase Shift=pi/4 Right end point=3pi/4 X-axis: yes Y-axis: No Vertical Shift: None?

zepdrix (zepdrix):

What does `right end point` mean? :o

OpenStudy (anonymous):

Right End Point is the ending of the complete cycle. Formula I was give is c/b + 2pi/b. In other words, the (phase shift)+(period).

zepdrix (zepdrix):

Ah oh ok ok ok :) End of one cycle, starting from x=0. Ok good! Looks great so far! And what does `x-axis` mean? Symmetry?

OpenStudy (anonymous):

X-axis orientation. In other words, is it reflected over the x-axis?

OpenStudy (anonymous):

Same for Y-axis.

zepdrix (zepdrix):

Oh oh I see, the negative tells you that it's reflected over the x-axis. Ok good good good. If you're listing the reflection separately, then you might want to list the amplitude as just the `magnitude` of the wave. Meaning drop the negative. It's amplified by 2, the negative just tells us about reflection. But you probably know your teacher better than I do, so I dunno :)

OpenStudy (anonymous):

I guess I wanted to make sure my math was right for the period and right end point. And wasn't completely sure of the vertical shift.

zepdrix (zepdrix):

Period is 2pi over B. 2pi divided by 4 simplifies to pi/2. good good good. pi/4 + pi/2 = 3pi/4 yay!

OpenStudy (anonymous):

She simply said that if \[a >0 then \it is \not reflected the x-axis. \]

OpenStudy (anonymous):

Oops. If a>0, then it does not reflect over the x-axis. If a<0, it does. Same concept for the y-axis only using b.

zepdrix (zepdrix):

Oh ok c: then she wants you to include it in the amplitude, ok cool

OpenStudy (anonymous):

So, for "Find the vertical shift", which I believe is represented by D, the answer would be none?

zepdrix (zepdrix):

Your function,\[\Large\rm f(x)=-2\sin(4x-\pi)\]is actually this,\[\Large\rm f(x)=-2\sin(4x-\pi)+0\]So yes, no vertical shift. That sounds correct :)

OpenStudy (anonymous):

Awesome! Thank you for the reassurance. End of the semester, could use all the correct answers possible.

zepdrix (zepdrix):

heh :3